ganeshie8 (ganeshie8):

\[1\le \pi(2^n)-\pi(2^{n-1}) \le \dfrac{2^n}{n-1}\]

ganeshie8 (ganeshie8):

\(\pi(x)\) is the prime counting function https://en.wikipedia.org/wiki/Prime-counting_function

OpenStudy (thomas5267):

So there is at least one prime between \(2^n\) and \(2^{n+1}\) and there can only be \(\dfrac{2^n}{n-1}\) many primes between the two powers of two?

ganeshie8 (ganeshie8):

Exactly!

OpenStudy (thomas5267):

This is horrible...

ganeshie8 (ganeshie8):

Use any theorems that you're familiar with...

OpenStudy (thomas5267):

It would be extremely uncomfortable to have no primes between two consecutive powers of two but I have no idea how to prove it.

ganeshie8 (ganeshie8):

I have simply used below postulate for the first inequality : https://en.wikipedia.org/wiki/Bertrand%27s_postulate

ganeshie8 (ganeshie8):

The lower bound trivially follows from the Bertrand's postulate.. I'm still working on the upper bound...

OpenStudy (thomas5267):

Bertrand's postulate is the first inequality you cheater!

ganeshie8 (ganeshie8):

Haha you have a good eye ;)

OpenStudy (thomas5267):

It follows trivially implies you still have to do trivial work using Bertrand's postulate but this is Bertrend's postulate! Cheater!

ganeshie8 (ganeshie8):

I can do some trivial work if you insist Bertrand's postulate give \(k \lt p \lt 2k\). Letting \(k=2^{n-1}\) gives the lower bound in the present problem :)

OpenStudy (dan815):

hm

OpenStudy (dan815):

\[p_{n+1}<2P_n\] this looks

OpenStudy (thomas5267):

Can't even use PNT to "morally" justify the upper bound. :( \[ \begin{align*} \pi(x)&\sim\frac{x}{\ln(x)}\\ \pi(2^n)-\pi(2^{n-1})&\sim\frac{2^n}{n\ln(2)}-\frac{2^{n-1}}{(n-1)\ln(2)}\\ &\sim\frac{(n-1)2^n-n2^{n-1}}{n(n-1)\ln(2)}\\ &\sim\frac{2^{n-1}(2(n-1)-n)}{n(n-1)\ln(2)}\\ &\sim\frac{2^{n-1}(n-2)}{n(n-1)\ln(2)}\\ \end{align*} \]

OpenStudy (dan815):

i thnk i have a certain way of building up the primes

ganeshie8 (ganeshie8):

\(\pi(x)\sim\frac{x}{\ln(x)} \implies \pi(2^n) - \pi(2^{n-1})\sim 0 \) since \(\sim\) is an asymptotic equivalence right ?

OpenStudy (thomas5267):

Yes it is asymptotic equivalence.

ganeshie8 (ganeshie8):

Oh it cannot be 0 as x/lnx is not constant..

ganeshie8 (ganeshie8):

nvm, your work looks sound...

OpenStudy (thomas5267):

It might be mathematically correct, but I don't see how it even justifies the upper bound.

ganeshie8 (ganeshie8):

@dan815 proving \(p_{n+1}\lt 2p_n\) isn't trivial... so im simply using bertrand's postulate for the lower bound

ganeshie8 (ganeshie8):

yeah we can't squeeze out an inequality out of asymptotic bound

OpenStudy (thomas5267):

\[ 2^n-2^{n-1}=2^{n-1} \] If the upper bound is true, than between two consecutive powers of two the proportion of primes is at most \(\dfrac{2}{n-1}\).

ganeshie8 (ganeshie8):

are you saying (number of integers) / (number of primes) = 2/(n-1) ?

OpenStudy (thomas5267):

Yes. I am just rephrasing the upper bound.

ganeshie8 (ganeshie8):

that looks very interesting... as \(n\to\infty\), the ratio approaches 0

ganeshie8 (ganeshie8):

\[\lim\limits_{n\to\infty} \dfrac{\pi(n)}{n}=0\]

ganeshie8 (ganeshie8):

That is really a very useful conclusion to make out of the upper bound !

OpenStudy (thomas5267):

From Wikipedia,\[\pi(x)<1.25506\frac{x}{\ln(x)}\] Not sure what can be made out of this.

ganeshie8 (ganeshie8):

Perhaps we can use that result at the end of : \(\pi(2^n) - \pi(2^{n-1}) \le \dfrac{2^n}{n-1}\\~\\ \implies \sum\limits_{n=1}^m \pi(2^n) - \pi(2^{n-1}) \le \sum\limits_{n=1}^m \dfrac{2^n}{n-1}\\~\\ \implies \pi(2^m) \le \sum\limits_{n=1}^m \dfrac{2^n}{n-1} \)

ganeshie8 (ganeshie8):

Not really sure.. im feeling all dyslexic today...

OpenStudy (thomas5267):

The denominator is n-1 so it won't work?

ganeshie8 (ganeshie8):

Ahh need to fix that

ganeshie8 (ganeshie8):

\(\pi(2^n) - \pi(2^{n-1}) \le \dfrac{2^n}{n-1}\\~\\ \implies \sum\limits_{n=2}^m \pi(2^n) - \pi(2^{n-1}) \le \sum\limits_{n=2}^m \dfrac{2^n}{n-1}\\~\\ \implies \pi(2^m) \le 1+ \sum\limits_{n=2}^m \dfrac{2^n}{n-1} \)

OpenStudy (thomas5267):

Is \(\dfrac{2^n}{n-1}\) the Taylor series coefficient of some function? I remember it is.

ganeshie8 (ganeshie8):

ln(1-x) ?

OpenStudy (thomas5267):

No it isn't ln(1-x). \(\dfrac{2^n}{n-1}\) seems to grow too fast for the series to converge.

OpenStudy (thomas5267):

I was thinking of replacing the right hand side of the inequality with a function that has the same Taylor series expansion. Does \(\displaystyle\sum_{n=2}^\infty\frac{2^n}{n-1}x^n\) converges for \(\displaystyle |x|<\frac{1}{2}\)?

OpenStudy (ikram002p):

...